# Conceptualization, Methodology, Software, Validation, Formal analysis, Data Curation, Investigation, Writing – original draft, Writing - Review & Editing, Visualization

### v1

@@ Abstract
As the integration of artificial intelligence technology with the medical field deepens, medication recommendation, as a crucial subfield, demonstrates significant potential value. Medication recommendation combines patient medical history with biomedical knowledge to assist doctors in determining medication combinations more accurately and safely. Existing approaches based on molecular knowledge overlook the atomic geometric structure of molecules, failing to capture the high-dimensional characteristics and intrinsic physical properties of medications, leading to structural confusion and the inability to extract useful substructures from individual patient visits. To address these limitations, we propose BiMoRec, which overcomes the inherent lack of molecular essential information in 2D molecular structures by incorporating 3D molecular structures and atomic properties. To retain the fast response required of recommendation systems, BiMoRec maximizes the mutual information between the two molecular modalities through bimodal graph contrastive learning, achieving the integration of 2D and 3D molecular graphs, and finally distills substructures through interaction with single patient visits. Specifically, we use deep learning networks to construct a pre-training method to obtain representations of 2D and 3D molecular structures and substructures, and we use contrastive learning to derive mutual information. Subsequently, we generate fused molecular representations through a trained GNN module, re-determining the relevance of substructure representations in conjunction with the patient’s clinical history information. Finally, we generate the final medication combination based on the extracted substructure sequences. Our implementation on the MIMIC-III and MIMIC-IV datasets demonstrates that our method achieves state-of-the-art performance. Compared to the next best baseline, our model improves accuracy by 1.8% while maintaining the same level of DDI as the baseline. Our source code is publicly available at: https://github.com/guangyunms/BiMoRec.
@@
We introduce a medication recommendation framework based on dual molecular modalities and patient-molecule interaction distillation, named BiMoRec.
For the first time in the field of medication recommendation, we incorporate dual molecular modalities, addressing the issue of structural confusion among molecules.
We propose a method that extracts multi-node key molecular substructures through single-visit interactions and performs molecular distillation using the latest visit, resolving the issue of substructure relevance bias.
@@ 1 Introduction
As medical demands continually increase and artificial intelligence technology progresses, smart healthcare becomes a robust support for addressing the future human challenges in medical resource distribution. Consequently, medical image processing, medical data analysis, and medication recommendation have emerged as significant directions for development. Medication recommendation, particularly, as it directly impacts medical outcomes, exhibits exceptionally high application value. Meanwhile, in the fields of biology and medicine, research activities such as molecular property prediction or molecular generation also propel deeper advancements in smart healthcare.
Medication recommendation [32, 3, 16] utilizes medical data analysis and deep learning technologies to recommend personalized medication combinations based on a patient’s clinical history and medication knowledge. Its distinction from traditional recommendations lies in the Drug-Drug Interactions (DDI) [2, 17, 6] that constrain the outcomes of medication recommendations, necessitating attention to safety. On the other hand, in traditional recommendation systems, the recommendations for current items primarily depend on the user’s interaction information with previous items [20, 5]. However, for medication recommendations, the relationships between a patient’s diagnostic information and medical entities play a more crucial role in the recommendation outcomes.
Early work on medication recommendation [27, 1, 28] focuses solely on the current patient visit, neglecting the longitudinal medication history of past consultations. Subsequent efforts, incorporating sequential recommendation approaches [14, 10, 35], achieve notable recommendation results. However, the issue of medication safety remains insufficiently addressed. Recent research concentrates on the impact of 2D molecular structures on medication recommendations [32, 33], making significant advancements. Nevertheless, this research overlooks the atomic geometric structure of molecules, leading to potential identical representations of molecules with similar 2D structures but different 3D structures. This structural confusion reduces the accuracy of medication recommendations.
@@
In this study, we identify two specific manifestations of structural confusion that lead to reduced accuracy. The first is that molecules with similar 2D structures but different 3D geometric structures are problematic because molecular reactions and bindings primarily occur in three-dimensional space. The limitations of 2D structures prevent us from accurately modeling the spatial relationships and interaction forces between molecules, thereby affecting the accuracy of medication recommendations. The second issue is that molecules segmented into smaller substructures do not represent the actual molecular states; they lose their original integrity and stability, and their 3D structures cannot be accurately reproduced in three-dimensional space. The molecular representations expressed by these substructures are also incomplete and inaccurate. Figure 1 shows the mapping relationships of molecular conformations, where even the same 2D molecular structures lead to different efficacies due to minor variations in 3D spatial configurations. We believe that by incorporating the atomic geometric configuration of molecules and completing the 3D information of substructures through comparative learning between bimolecular modalities, we can enhance the overall capability of medication recommendation.
@@
Following this line, we design the BiMoRec model, which addresses the confusion of 2D molecular structures and the inability of substructures to represent true molecular states. Initially, we devise a bimodal graph comparative learning phase, where the bimodal data of molecules serve as input. The 2D modality, through the GIN network, obtains low-dimensional structural embeddings of the molecules, while for the 3D modality, we first acquire each molecule’s 3D data structure using the SDF description file of the molecule, and simultaneously complete the 3D description for the segmented substructures. Subsequently, through the atomic geometric vector machine (GVP) [11], we obtain high-dimensional structural embeddings of the molecules. We then compare the mutual information between the two modal graphs [4], integrating the information of both modalities, allowing the 2D encoder to learn the high-dimensional structural representations of 3D. After completing the comparative learning, we proceed to study the correlation between a given patient’s health condition and the bimolecular granular structure, and finally, by capturing the superficial correlations of medications through the patient’s medication history, we correct any deviations in the recommendation results. Through cross-layer synergistic enhancement, this process yields more accurate and safer medication combinations.
@@
Compared to the latest methods, our model’s introduction of the 3D molecular modality not only completes the molecular structural representation but also employs a pre-trained module with parameters learned in advance. This does not affect the performance of subsequent stages, thereby achieving a balance between accuracy and speed, effectively reducing time costs. Moreover, this module is highly extensible and can be integrated with any medication recommendation framework that utilizes molecular knowledge.
@@
  • We introduce the high-dimensional 3D structure of molecules and complete the 3D information of substructures. By using bimodal graph contrastive learning to integrate the knowledge of both modalities, it effectively corrects the structural confusion problem and optimizes the representation space of molecules.
  • To improve accuracy and safety, we obtain key substructure sequences through single interactions between patient and molecule, and perform molecular distillation through the latest visit to derive the recommended medication combinations. This approach ensures safety by leveraging both molecular knowledge and historical information.
@@
Below, we provide an overview of the different sections of the paper: (1) Introduction: Introduces the main innovations and the motivation behind this work. (2) Related Work: Summarizes typical studies and current hotspots in medication recommendation fields. (3) Methods: Presents the core ideas of the model and the specific details of its implementation. (4) Experiments: Describes the experimental background, baselines, testing methods, and specific experimental results. (5) Discussion: Delves into an analysis of the experimental results and proposes a series of supportive experiments. (6) Conclusion: Summarizes the research findings and offers perspectives on future research directions.
@@ 2.1 Medication Recommendation
The field of medication recommendation has developed rapidly in recent years, with its primary objective being to provide the most suitable medication combination based on the patient’s specific condition.
Early works in medication recommendation like [34] are instance-based, prescribing on single-visit data, and viewing medication recommendation as a multi-instance and multi-label classification task. After that, studies like [7] employ sequential models to integrate longitudinal visit data. Recently, the research falls into three categories. The first category of work, exemplified by studies like [15, 36, 16], treats medication recommendation as a package, sequence, and other mature recommendation tasks, utilizing advanced algorithms to enhance model quality. The second category of work integrates techniques from other fields, such as [31, 29, 37]. They combine the task of medication recommendation with translation models, generative models, or technologies from other domains, achieving a method that transforms patient states into medication combinations. The third category of work, such as [23, 32, 33, 3], incorporates external knowledge related to medications to fill the knowledge gaps in the data, thereby achieving more accurate representations of medications.
@@ 3.1 Medical Entity
This article introduces the concept of medical entities, which represent relatively comprehensive medical concepts in medical data. In this study, medical entities primarily include three types: diseases, surgeries, and medications, denoted as $\mathcal{D}=\{d_{1},d_{2},\ldots\}$, $\mathcal{P}=\{p_{1},p_{2},\ldots\}$, and =$\mathcal{M}=\{m_{1},m_{2},\ldots\}$, respectively. Additionally, molecular modal data is derived from external knowledge and thus does not belong to medical entities.
@@ 3.2 Input and Output
The model utilizes Electronic Health Record (EHR) as the data source, encompassing extensive records of patient visits and treatments. Each patient record is represented as $\mathcal{H}$, containing multiple visit records =$\mathcal{H}=\{v_{1},v_{2},\ldots,v_{t}\}$, where $v_{t}$ corresponds to the clinical information related to the $t_{th}$ visit. Each visit includes three elements: $\mathcal{D}_{t}\in\{0,1\}^{|\mathcal{D}|},\mathcal{P}_{t}\in\{0,1\}^{\mathcal{P}|},\mathcal{D}_{t}\in\{0,1\}^{|\mathcal{M}|}$, all of which are represented using multi-hot encoding. The output of the model is denoted as ^$\hat{\mathcal{M}}_{t}$, which is the predicted medication combination for the visit $v_{t}$.
@@ 3.3 DDI Matrix
Drug-Drug interactions are a focal point in the field of medication recommendation, as they indicate that certain drug combinations may pose validated safety risks to patients. By excluding drug combinations with safety risks, adverse events caused by such interactions can be reduced, thereby enhancing the safety of drug therapy. The real-world DDI rate is about 0.06. Our DDI information is extracted from the Adverse Event Reporting Systems [9]. We represent DDI information using a binary matrix ${\mathbf{M}^{ddi}}\in\{0,1\}^{|\mathcal{M}|\times|\mathcal{M}|}$, where $\mathbf{M}^{ddi}_{ij}=1$ indicates the presence of an interaction between $m_{i}$ and $m_{j}$. A high frequency of DDI suggests potential safety issues in recommended results.
@@ 4 Methods
Figure 2: The bottom left of the diagram shows the pre-training and integration process of the bimodal molecular structures, the results of which are applied to the downstream tasks depicted in the top left. The top left section explains how molecular knowledge and patient representations are integrated across layers. This is done to recommend medication combinations at the granularity of single visits, thereby achieving deep knowledge interaction and distilling and extracting molecular substructures at each time point.The right side of the diagram models the sequential relationships of the interaction results and maps them to recommendation probabilities, ultimately yielding the final recommendation results.
@@
(2) In the molecular multi-distillation stage, our objective is to derive single-visit representations of the patient from clinical records. Using the encoders obtained during the pre-training phase, we encode the molecules and their substructures of the medications. Through the collaborative enhancement of deep knowledge and shallow representations, we obtain medication combination results across multiple visits, and based on the current patient’s status, we acquire medication recommendation results under different temporal contexts.
(3) In the stage of recommendation path fusion, our goal is to integrate the sequential relationships of medication recommendations at different time points to obtain the sequence of multiple recommendation results. Subsequently, based on the results of the fusion of multiple time-series environments, we adjust the probability bias for each medication according to the patient’s current health condition, thereby recommending medications that exceed a predetermined threshold.
@@ 4.1 Pre-training
In terms of embedding representations, 2D modal data utilizes Graph Neural Networks (GNN) to encode the molecular structure of drugs and obtain holistic-level representations. GNN iteratively updates node features by aggregating features of neighboring nodes, i.e., updating representations of atomic nodes through specific aggregation strategies. In general, we have:
@@
where $\mathbf{f}_{n}^{(k)}(u)$ is the node feature of node $u$ in the $k$-th layer, $\mathbf{f}_{n}^{(0)}(u)$ is the initial feature of node $u$, $\mathbf{f}_{e}(u,v)$ is the edge feature of edge $(u,v)$, and $\mathcal{N}(u)$ denotes the neighbor of node $u$ on the graph.
A drug molecule is composed of atoms and bonds, and can be represented as a graph $G=(V,E)$, where $V$ is the set of atomic nodes in the molecular graph, and $E$ is the set of chemical bond edges in the molecular graph. Given a molecule $G$, the representation of the entire graph is generated by $L$ layers of GNN, and then aggregated by a readout function to combine all node features in the $L$th layer,
@@
In this paper, we use GIN [30] as the encoder for drug molecular graphs, which is commonly used in recent work related to drug recommendation and molecular research. We utilize pre-trained encoders as the foundation for subsequent tasks, and the overall-level representations they output are employed for downstream tasks. All drug molecular representations constitute an embedding table $\mathbf{E}_{m}\in R^{|\mathcal{M}|\times h}$, where each row represents a molecular’s representation.
@@
BiMoRec also integrates the 3D modal structure of molecules to enhance molecular representations. Similarly, given the 3D coordinates of an atom, we represent the molecular structure as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the nodes $\mathcal{V}$ constitute the set of atoms in the molecule, and we define an edge $\mathcal{E}$ if the distance between atoms is less than 4.5 Å. We directly utilize the 3D coordinates of atoms or edges as node features and edge features. In graph $\mathcal{G}$, we associate each node and edge with geometric features such as 3D coordinates, as well as scalar features describing chemical properties. Thus, we employ dedicated Geometric Vector Perceptrons (GVPs) to construct the molecular structure. Its main advantage lies in its design specifically considering 3D data, enabling direct learning of structural representations from raw atomic coordinates without the need to construct features that are invariant to rotation and translation.
@@
In model construction, we use the tuple $\mathbf{v}_{i}=(\mathbf{v}^{v}_{i},\mathbf{v}^{s}_{i})$ to represent the vector features and scalar features of atom $i$ separately, where $\mathbf{v}^{v}_{i}\in\mathbb{R}^{\mu\times 3}$ is a list of vector features in $\mathbb{R}^{3}$, and $\mathbf{v}^{s}_{i}\in\mathbb{R}^{\nu}$ is a list of scalar features ($\mu$ and $\nu$ are the number of vector features and scalar features, respectively). The edge feature $\mathbf{e}_{ij}=(\mathbf{e}^{v}_{ij},\mathbf{e}^{s}_{ij})$ of edge $(i,j)$ has a similar meaning. The molecular GVP transforms node and edge features through $L$ graph convolution layers to obtain the representation of the input molecule. Specifically, in the $i$th layer, each node aggregates embeddings from neighboring nodes and edges, and then updates its own representation:
$\mathbf{h}_{i}^{(\ell)}=\mathbf{h}_{i}^{(\ell-1)}+g\Bigg{(}\mathbf{h}_{i}^{(\ell-1)}+\frac{1}{|\mathcal{N}(i)|}\sum_{j\in\mathcal{N}(i)}\mathbf{m}_{ji}^{(\ell)}\Bigg{)},$ (4)
@@
In this context, $g(\cdot)$ denotes a sequence of three GVP layers, and $\mathcal{N}(i)$ represents the set of neighboring nodes of node $i$ in $\mathcal{G}$. The embedding of node $i$ at layer $\ell$ is $\mathbf{h}^{(\ell)}_{i}$ (specifically, at layer 0, $\mathbf{h}^{(0)}_{i}=\mathbf{v}_{i}$). The message from node $j$ to node $i$ is $\mathbf{m}^{(\ell)}_{ji}$, computed using another sequence of GVP layers: $\mathbf{m}^{(\ell)}_{ji}=g(\text{concat}(\mathbf{h}^{(\ell-1)}_{j},e_{ji}))$, where $\text{concat}(\cdot)$ denotes the concatenation operation of two embeddings. After the final layer of molecular GVP, we also apply a global additive pooling operation to aggregate all node representations into the scalar representation $\mathbf{h}_{d}\in\mathbb{R}^{128}$ of the input drug, obtaining the 3D embedding matrix $\mathbf{E}_{mt}$ of the molecule, where each row represents a molecule’s representation.
@@
Contrastive learning achieves two objectives. The first objective is to make molecules with the same index $i$ increasingly similar. The second objective is to enforce differences between negative samples $\mathbf{E}^{(i)}_{a}$ and $\mathbf{E}^{(k)}_{b}$, where $i\neq k$. If the 2D and 3D representations in a batch come from different molecules, their representations should differ. We use the currently popular NTXent loss to jointly optimize our model:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\left[\log\frac{e^{sim(\mathbf{E}^{(i)}_{a},\mathbf{E}^{(i)}_{b})/\tau}}{\sum_{k\neq i}^{N}e^{sim(\mathbf{E}^{(i)}_{a},\mathbf{E}^{(k)}_{b})/\tau}}\right]$ (5)

### v2

@@ Abstract
As the integration of artificial intelligence technology with the medical field deepens, medication recommendation, as an important subfield, demonstrates immense potential value. Medication recommendation combines patient medical history with biomedical knowledge[31, 42, 43] to assist doctors in determining medication combinations more accurately and safely. Existing works based on molecular knowledge neglect the 3D geometric structure of molecules and fail to learn the high-dimensional information of medications, leading to structural confusion. Additionally, it does not extract key substructures from a single patient visit, resulting in the failure to identify medication molecules suitable for the current patient visit. To address the above limitations, we propose a bimodal molecular recommendation framework named BiMoRec, which introduces 3D molecular structures to obtain atomic 3D coordinates and edge indices, overcoming the inherent lack of high-dimensional molecular information in 2D molecular structures. To retain the fast training and prediction efficiency of the recommendation system, we use bimodal graph contrastive pretraining to maximize the mutual information between the two molecular modalities, achieving the fusion of 2D and 3D molecular graphs and re-evaluating substructures at the visit level. Specifically, we use deep learning networks to construct a pretraining method that acquires 2D and 3D molecular structure representations and substructure representations, and obtain mutual information through contrastive learning. We then generate fused molecular representations using the trained GNN module and re-determine the relevance of substructure representations in combination with the patient’s clinical history. Finally, we generate the final medication combination based on the extracted substructure sequences. Our implementation on the MIMIC-III and MIMIC-IV datasets demonstrates that our method achieves state-of-the-art performance. Compared to the second-best baseline, our model improves accuracy by 2.07%, with DDI at the same level as the baseline. Our source code is publicly available at: https://github.com/guangyunms/BiMoRec.
@@
We introduce a medication recommendation framework based on bimolecular modalities and dynamic enhancement of patient-molecule interactions, named BiMoRec.
We introduce molecular bimodality into the field of medication recommendation for the first time, addressing the issue of structural confusion between molecules.
We propose a method for molecular enhancement and evaluation at the single-visit granularity, extracting key molecular substructures across multiple time points, and re-enhancing them based on the latest visit. This approach addresses the issue of substructure correlation bias.
@@ 1 Introduction
Medication recommendation systems are an important research field in modern medical informatics, aiming to provide personalized medication suggestions to doctors and patients by analyzing patients’ personal health data (e.g., electronic health records (EHRs)) and medical literature. With the rapid advancements in big data, artificial intelligence, and machine learning technologies, smart healthcare has gradually become an important means of addressing future challenges in medical resource distribution. Unlike traditional medication recommendations, which primarily rely on doctors’ experience and guidelines, modern medication recommendation systems can utilize vast clinical datasets and literature databases, automatically extracting useful information and performing intelligent analysis, thereby improving the accuracy and efficiency of recommendations.
Unlike recommendations in other fields (such as e-commerce platforms), medication recommendation not only focuses on the accuracy of the recommendations but also emphasizes their safety. This is because different medications can interact with each other, potentially causing adverse effects and impacting the patient’s health. For instance, both aspirin and warfarin are anticoagulant medications, but their simultaneous use may increase the risk of bleeding. A study [2] shows that in the largest publicly available benchmark dataset, MIMIC-III, patients have an average of 20 different medications. Therefore, eliminating all potential drug-drug interaction (DDI) [3, 23, 6] remains a significant challenge.
Early medication recommendation work focuses on the patient’s latest visit [36, 1, 38], considering the current state as a crucial factor for medication recommendation. Subsequent researchers incorporate the patient’s longitudinal visit records [19, 12, 45], achieving better results. However, another important indicator for medication recommendation, safety, remains inadequately addressed. To balance accuracy and safety, recent research focuses on the impact of 2D molecular structures on medication recommendation [42, 43], making significant progress. Representing medications as finer-grained molecular substructures helps in modeling the pharmacological properties of medications and DDI, significantly improving model accuracy and safety, ushering medication recommendation models into a new stage. However, this research overlooks the geometric structure of atoms, causing molecules with similar 2D structures but different 3D structures to potentially have the same representation. This structural confusion reduces the accuracy of medication recommendation. The high-dimensional information contained in 3D structures (such as 3D atomic coordinates and their edges) can more clearly distinguish different molecular structures, optimizing the molecular representation space to address this issue.
@@
In this study, we identify two specific manifestations of structural confusion that lead to reduced accuracy. The first is that molecules with similar 2D structures but different 3D geometric structures are problematic because molecular reactions and bindings primarily occur in three-dimensional space. The limitations of 2D structures prevent us from accurately modeling the spatial relationships and interaction forces between molecules, thereby affecting the accuracy of medication recommendations. The second issue is that molecules segmented into smaller substructures do not represent the actual molecular states; they lose their original integrity and stability, and their 3D structures cannot be accurately reproduced in three-dimensional space. The molecular representations expressed by these substructures are also incomplete and inaccurate. Figure 1 shows the mapping relationships of molecular conformations, where even the same 2D molecular structures lead to different efficacies due to minor variations in 3D spatial configurations. We believe that by incorporating the 3D geometric configuration of molecules and completing the 3D information of substructures through comparative learning between bimolecular modalities, we can enhance the overall capability of medication recommendation.
@@
Following this line, we design the BiMoRec model, which addresses the confusion of 2D molecular structures and the inability of substructures to represent true molecular states. Initially, we devise a bimodal graph comparative learning phase, where the bimodal data of molecules serve as input. The 2D modality, through the GIN network [40], obtains low-dimensional structural embeddings of the molecules, while for the 3D modality, we first obtain the molecular SDF description files from the molecules’ SMILES strings, acquiring the corresponding 3D data for each molecule, and simultaneously complete the 3D description for the segmented substructures. Subsequently, through the 3D geometric vector machine (GVP) [13, 25], we obtain high-dimensional structural embeddings of the molecules. We then compare the mutual information between the two modal graphs [5], integrating the information of both modalities, allowing the 2D encoder to learn the high-dimensional structural representations of 3D. After the contrastive learning phase, we enhance the set of molecules used within a single visit by learning the correlation between the given patient’s health status and the bimodal molecular structures. We capture the highly correlated substructures for each visit using the patient’s multiple visits and further enhance the recommendation results using the latest visit. Through multi-granularity collaborative enhancement, we obtain more accurate and safe medication combinations.
@@
Compared to the latest methods, our model introduces a 3D molecular modality that not only completes the structural representation of molecules but also uses a pre-training module with pre-learned parameters, ensuring that the model performance in subsequent stages is not affected. Our pre-training module conducts bimodal molecular contrastive learning before formal training, meeting the established objectives while maintaining the expected model efficiency. Moreover, this module is highly extensible and can be integrated with any medication recommendation framework that utilizes molecular knowledge.
@@
  • We introduce the high-dimensional 3D structure of molecules for the first time in medication recommendations and complete the 3D information of substructures. By using bimodal graph contrastive learning to integrate the knowledge of both modalities, it effectively corrects the structural confusion problem and optimizes the representation space of molecules.
  • This paper enhances molecular embeddings based on the visit-level representations of patients and re-evaluates the set of substructures to obtain key substructure sequences for each visit. It further enhances the molecular embeddings based on the patient’s latest visit to emphasize molecular features that are more relevant to the current patient condition, improving model performance while ensuring a balance between accuracy and safety.
@@
Below, we provide an overview of the different sections of the paper: (1) Introduction: Introduces the main innovations and the motivation behind this work. (2) Related Work: Summarizes typical studies and current hotspots in medication recommendation fields. (3) Problem Definition: Introduces the medical concepts, entities, inputs, and outputs used. (4) Methods: Presents the core ideas of the model and the specific details of its implementation. (5) Experiments: Describes the experimental background, baselines, testing methods, and specific experimental results. (6) Discussion: Delves into an analysis of the experimental results and proposes a series of supportive experiments. (7) Conclusion: Summarizes the research findings and offers perspectives on future research directions.
@@ 2.1 Medication Recommendation
In recent years, the field of medication recommendation has undergone multiple development stages, aiming to provide personalized medication combinations for patients.
Early works in medication recommendation include instance-based methods [44], which prescribe medications based on single-visit data and view medication combination recommendations as multi-instance and multi-label classification tasks. Subsequent works introduce sequential models [7] to model longitudinal data.
Recent medication recommendation works are categorized into three types. The first type views medication recommendation as a package, sequence, and other mature recommendation tasks [20, 46, 22], utilizing advanced algorithmic models to improve recommendation quality. The second type combines the medication recommendation task with translation models [39] and generative models [41, 47] to derive medication combinations based on the patient’s personal visit information. The last type utilizes relevant external knowledge to fill in missing information in the data, thereby improving model performance, as in [31, 42, 43, 4].
@@ 3.1 Medical Entity
This article introduces the concept of medical entities, which represent relatively comprehensive medical concepts in medical data. In this study, medical entities primarily include three types: diseases, surgeries, and medications, denoted as $\mathcal{D}=\{d_{1},d_{2},\ldots\}$, $\mathcal{P}=\{p_{1},p_{2},\ldots\}$, and =$\mathcal{M}=\{m_{1},m_{2},\ldots\}$. Additionally, molecular modal data is derived from external knowledge and thus does not belong to medical entities.
@@ 3.2 Input and Output
The model utilizes Electronic Health Record (EHR) as the data source, encompassing extensive records of patient visits and treatments. Each patient record is represented as $\mathcal{H}$, containing multiple visit records =$\mathcal{H}=\{v_{1},v_{2},\ldots,v_{t}\}$, where $v_{t}$ corresponds to the clinical information related to the $t_{th}$ visit. Each visit includes three elements: $\mathcal{D}_{t}\in\{0,1\}^{|\mathcal{D}|},\mathcal{P}_{t}\in\{0,1\}^{|\mathcal{P}|},\mathcal{D}_{t}\in\{0,1\}^{|\mathcal{M}|}$, all of which are represented using multi-hot encoding. The output of the model is denoted as ^$\hat{\mathcal{M}}_{t}$, which is the predicted medication combination for the visit $v_{t}$.
@@ 3.3 DDI Matrix
Drug-Drug interactions are a focal point in the field of medication recommendation, as they indicate that certain medication combinations may pose validated safety risks to patients. By excluding medication combinations with safety risks, adverse events caused by such interactions can be reduced, thereby enhancing the safety of medication therapy. The real-world DDI rate is about 0.06. Our DDI information is extracted from the Adverse Event Reporting Systems [10]. We represent DDI information using a binary matrix ${\mathbf{M}^{ddi}}\in\{0,1\}^{|\mathcal{M}|\times|\mathcal{M}|}$, where $\mathbf{M}^{ddi}_{ij}=1$ indicates the presence of an interaction between $m_{i}$ and $m_{j}$. A high frequency of DDI suggests potential safety issues in recommended results.
@@ 4 Methods
Figure 2: The pre-training phase illustrated in the bottom left side of the figure demonstrates the pre-training fusion process of bimodal molecular structures, and its results are applied to the downstream tasks shown in the top left side. The molecular multi-enhancement phase in the top left side explains how molecular knowledge interacts with patient representations at the visit level, dynamically enhancing substructures at the granularity of a single visit. The trajectory fusion phase on the right side of the figure models the sequence relationships of the interaction results and maps them to recommendation probabilities, obtaining the final recommendation results.
@@
(2) In the molecular multiple enhancement phase, our goal is to obtain the patient’s single-visit representation from clinical records and encode the medication molecules and their substructures using the encoder obtained from the pre-training phase. By interacting molecular knowledge with the patient representation, we enhance and re-evaluate the representations of the molecules and their substructures. Finally, we further enhance the combinations of substructures at different time points based on the patient’s current visit status.
(3) In the medication evolution trajectory fusion stage, our objective is to integrate the enhanced substructure representations from different time sequences to obtain the sequential relationships of multiple recommendation results. Subsequently, based on the results of the multi-sequence environment integration, we adjust the probability bias of each medication according to the patient’s current health condition, thereby recommending medications that exceed the preset threshold.
@@ 4.1 Pre-training
In terms of embedding representations, 2D modal data utilizes Graph Neural Networks (GNN) [40] to encode the molecular structure of drugs and obtain holistic-level representations. GNN iteratively updates node features by aggregating features of neighboring nodes, i.e., updating representations of atomic nodes through specific aggregation strategies. In general, we have:
@@
where $\mathbf{f}_{n}^{(k)}(u)$ is the node feature of node $u$ in the $k$-th layer, $\mathbf{f}_{n}^{(0)}(u)$ is the initial feature of node $u$, $\mathbf{f}_{e}(u,v)$ is the edge feature of edge $(u,v)$, and $\mathcal{N}(u)$ denotes the neighbor of node $u$ on the graph. $\mathrm{COMBINE}$ is responsible for fusing the features of the neighborhood with the features of the central node, thereby updating the features of the central node.
A medication molecule is composed of atoms and bonds, and can be represented as a graph $G=(V,E)$, where $V$ is the set of atomic nodes in the molecular graph, and $E$ is the set of chemical bond edges in the molecular graph. Given a molecule $G$, the representation of the entire graph is generated by $L$ layers of GNN, and then aggregated by a readout function to combine all node features in the $L$th layer,
@@
In this paper, we use GIN as the encoder for drug molecular graphs, which is commonly used in recent work related to drug recommendation and molecular research [11, 27, 37]. We utilize pre-trained encoders as the foundation for subsequent tasks, and the overall-level representations they output are employed for downstream tasks. All drug molecular representations constitute an embedding table $\mathbf{E}_{m}\in R^{|\mathcal{M}|\times h}$, where each row represents a molecular’s representation.
@@
BiMoRec also integrates the 3D modal structure of molecules to enhance molecular representations. Similarly, given the 3D coordinates of an atom, we represent the molecular structure as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the nodes $\mathcal{V}$ constitute the set of atoms in the molecule, and we define an edge $\mathcal{E}$ if the distance between atoms is less than 4.5 Å. We directly utilize the 3D coordinates of atoms or edges as node features and edge features. In graph $\mathcal{G}$, we associate each node and edge with geometric features such as 3D coordinates, as well as scalar features describing chemical properties. Thus, we employ dedicated Geometric Vector Perceptrons (GVPs) to construct the molecular structure. Its main advantage lies in its design specifically considering 3D data, enabling direct learning of structural representations from raw atomic coordinates without the need to construct features that are invariant to rotation and translation. The equivariance and invariance of GVP are reflected in that, if we have $\text{GVP}(\mathbf{s},\mathbf{V})=(\mathbf{s^{\prime}},\mathbf{V^{\prime}})$, then for any combination of rotations and reflections $R$, we have $\text{GVP}(\mathbf{s},R(\mathbf{V}))=(\mathbf{s^{\prime}},R(\mathbf{V^{\prime}}))$.
For the feature learning process of 3D information in the GVP neural network layer, given $\mathbf{f}=(\mathbf{s},\mathbf{V})$, where $\mathbf{s}\in\mathbb{R}^{\nu}$ is a list of scalar features and $\mathbf{V}\in\mathbb{R}^{\mu\times 3}$ is a list of vector features in $\mathbb{R}^{3}$ ($\mu$ and $\nu$ are the number of vector features and scalar features, respectively). The specific feature learning process is as follows:
$\mathbf{s}_{hv}=\text{CONCAT}({||\mathbf{W}_{h}V||}_{2},\mathbf{s}),$ (4)
$\mathbf{s^{\prime}}=\sigma(\mathbf{W}_{v}\mathbf{s}_{hv}+b),$ (5)
$\boldsymbol{v}_{\mu}={||\mathbf{W}_{\mu}\mathbf{W}_{h}V||}_{2},$ (6)
$\mathbf{V^{\prime}}=\sigma(\boldsymbol{v}_{\mu})\odot\boldsymbol{v}_{\mu}$ (7)
where $W_{h}$, $W_{v}$, and $W_{\mu}$ are three weight matrices for transformation, $b$ is a bias term, and $\odot$ represents element-wise multiplication.
@@
In model construction, we use the tuple $\mathbf{f}_{i}=(\mathbf{s}_{i},\mathbf{V}_{i})$ to represent the vector features and scalar features of atom $i$ separately, where $\mathbf{V}_{i}\in\mathbb{R}^{\mu\times 3}$ is a list of vector features in $\mathbb{R}^{3}$, and $\mathbf{s}_{i}\in\mathbb{R}^{\nu}$ is a list of scalar features. The edge feature $\mathbf{e}_{ij}=(\mathbf{e}^{v}_{ij},\mathbf{e}^{s}_{ij})$ of edge $(i,j)$ has a similar meaning. The molecular GVP transforms node and edge features through $L$ graph convolution layers to obtain the representation of the input molecule. Specifically, in the $i$ layer, each node aggregates embeddings from neighboring nodes and edges, and then updates its own representation:
$\mathbf{h}_{i}^{(l)}=\mathbf{h}_{i}^{(l-1)}+g\Bigg{(}\mathbf{h}_{i}^{(l-1)}+\frac{1}{|\mathcal{N}(i)|}\sum_{j\in\mathcal{N}(i)}\mathbf{m}_{ji}^{(l)}\Bigg{)},$ (8)
@@
In this context, $g(\cdot)$ denotes a sequence of three GVP layers, and $\mathcal{N}(i)$ represents the set of neighboring nodes of node $i$ in $\mathcal{G}$. The embedding of node $i$ at layer $l$ is $\mathbf{h}^{(l)}_{i}$ (specifically, at layer 0, $\mathbf{h}^{(0)}_{i}=\mathbf{f}_{i}$). The message from node $j$ to node $i$ is $\mathbf{m}^{(l)}_{ji}$, computed using another sequence of GVP layers: $\mathbf{m}^{(l)}_{ji}=g(\text{concat}(\mathbf{h}^{(l-1)}_{j},e_{ji}))$, where $\text{concat}(\cdot)$ denotes the concatenation operation of two embeddings.
After the final layer of molecular GVP, we also apply a global additive pooling operation to aggregate all node representations into the scalar representation $\mathbf{h}_{d}\in\mathbb{R}^{128}$ of the input drug, obtaining the 3D embedding matrix $\mathbf{E}_{mt}$ of the molecule, where each row represents a molecule’s representation.
@@
Contrastive learning achieves two objectives. The first objective is to make molecules with the same index $i$ increasingly similar. The second objective is to enforce differences between negative samples $\mathbf{E}_{m_{i}}$ and $\mathbf{E}_{n_{k}}$, where $i\neq k$, i and k respectively represent the $i_{th}$ embedding in the 2D embedding table and the $k_{th}$ embedding in the 3D embedding table. If the 2D and 3D representations in a batch come from different molecules, their representations should differ. We use the currently popular NTXent loss [25, 5] to jointly optimize our model:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\left[\log\frac{e^{sim(\mathbf{E}_{m_{i}},\mathbf{E}_{n_{i}})/\tau}}{\sum_{k\neq i}^{N}e^{sim(\mathbf{E}_{m_{i}},\mathbf{E}_{n_{k}})/\tau}}\right]$ (9)